OpenStudy (anonymous):

please help me Kent has two similar cylindrical pipes, Pipe A and Pipe B. The radius of Pipe A is 6 cm, and the radius of Pipe B is 2 cm. What is the ratio of the volume of Pipe A to the volume of Pipe B? (1 points) 3:1 9:1 27:1 81:1

OpenStudy (anonymous):

oh i think its A because 6/2=3 and 2/2=1 which = ratio 3:1

OpenStudy (anonymous):

try 9/1 vol=pi*h*r^2 A=pi*h*6*6=36 B=pi*h*2*2=4............ratio 36/4=9/1

OpenStudy (anonymous):

and where did u get all of that?

OpenStudy (anonymous):

@mathstudent55 could you help me out here a bit

OpenStudy (anonymous):

that's the standard volume of a cylinder..... area of the circle * height area of circle is pi*r^2 when i forget i just google for the formula this site is good for answering particular questions but many other sites will cover the topic better for instance purplemath.com will explain how to find volumes etc.